OpenStudy (anonymous):

Evaluate -4(a+b) - 10a/b when a=8 and b=5 Answer: a. -15 b. 36 c. -53 d. 625/8 e. -68

OpenStudy (anonymous):

-68 is ans

OpenStudy (mayankdevnani):

\[\Huge{\bf{\color{blue}{Many}\color{red}{Problems!!!}}}\]

OpenStudy (anonymous):

will you break it down for me so i can put it in my notes

OpenStudy (anonymous):

i dont have problems bro

OpenStudy (mayankdevnani):

ok!! sis

OpenStudy (mayankdevnani):

\[-4(8+5)-10\times \frac{8}{5}\]

OpenStudy (mayankdevnani):

is it your question? @skittles223

OpenStudy (anonymous):

hey when an x is in front of the 2 what is the x for

OpenStudy (mayankdevnani):

question is right?

OpenStudy (anonymous):

-4(8+5)-16(8/5)=0 ,-4(13)-16=0,-52-16=0,-68=0

OpenStudy (anonymous):

what is that i dont get that equation where did you get all the numbers from

OpenStudy (anonymous):

*16=10 made a mistake

OpenStudy (mayankdevnani):

\[-4\times13-2\times8\] \[-52-16=-68\]

OpenStudy (mayankdevnani):

ok @skittles223

OpenStudy (mayankdevnani):

right @thivitaa

OpenStudy (anonymous):

this is right

OpenStudy (mayankdevnani):

thnx.....

OpenStudy (anonymous):

okk ..

OpenStudy (anonymous):

question is so easy

OpenStudy (mayankdevnani):

yaa

OpenStudy (anonymous):

is it all the 16 suppose to be 10

OpenStudy (anonymous):

?

OpenStudy (anonymous):

wht question?

OpenStudy (anonymous):

is the 16's all suppose to be 10's

OpenStudy (anonymous):

-388/5 this is ans.

OpenStudy (anonymous):

to what

OpenStudy (anonymous):

the highest they got is 625/8

OpenStudy (anonymous):

wht is

OpenStudy (anonymous):

are we talking about this question Evaluate -4(a+b) - 10a/b when a=8 and b=5

OpenStudy (anonymous):

plz again write the question which r u want and which r u changes in q?

OpenStudy (anonymous):

Evaluate -4(a+b) - 10a/b when a=8 and b=5

OpenStudy (anonymous):

-4(8+5)-10x8/5=-68

OpenStudy (anonymous):

If a 30 foot flagpole casts a12 foot shadow, how tall is a mailbox casting a 18 inch shadow?

OpenStudy (anonymous):

will you do it step by step so i know how to do it next time im takin notes

OpenStudy (nubeer):

1 feet = 12 inches (pre - known) 18 inches = 1.5 feets. 30 feet = 12 feet shadow 30/12feet = 1 feet shadow 1.5 (30/12) = 1.5 feet shadow 3.75 foots tall mail box.

OpenStudy (anonymous):

@nubeer u r a teacher?

OpenStudy (nubeer):

nop.. why?

OpenStudy (anonymous):

@nubeer like how do you find the answer to that problem though

OpenStudy (anonymous):

they dont have 3.75

OpenStudy (anonymous):

why don't you convert to inches all the way... 30 ft = 360 inches , 12 ft is 144 inches , 144/360 = .4 18 inches times .4 = 7.2 inches thats how long it would be

OpenStudy (anonymous):

360/144 = 2.5 18 x 2.5 =45 inches ..sorry

OpenStudy (anonymous):

thanks spectrum i really appreciate it

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

this is 45 inch tall mailbox

OpenStudy (anonymous):

@nubeer i don't at all mean to be disrespectful but you post including the "steps" was kind of confusing to me.

OpenStudy (anonymous):

45 inches is 3.75 ft btw everyone..

OpenStudy (nubeer):

hmm yes maybe.. it's basically ratio method question . so was just trying to convey how normally the ratio method is done. obviously there are more many ways to solve.. but it's fine as long as its done right.

OpenStudy (anonymous):

It takes Karina 3.5 hours to drive from Davenport to Hogsdale and then 2 hours and 10 minutes to drive from Hogsdale to Paqua. On the route that she took, it is 720 miles from Davenport to Paqua. What was her average driving speed for the trip?

OpenStudy (hba):

Do You Know What Is The Average Speed Formula ?

OpenStudy (hba):

@skittles223

OpenStudy (anonymous):

no srry

OpenStudy (anonymous):

?

OpenStudy (hba):

Average speed = s / t

OpenStudy (anonymous):

what is s/t for and what are there numbers

OpenStudy (hba):

S=720 miles t=3.5 hrs + 2 hrs + 10 minutes= 5 hours and 40 minutes = 5 2/3 hours = 17/3 hours S is the distance t is the total time which we calculated by adding now put it in the formula :)

OpenStudy (hba):

And Yeah Don't Forget The Unit,You Can Loose Valuable Marks :) Got it ? @skittles223

OpenStudy (anonymous):

hba i tried to get the answer and i cant dont know how to do

OpenStudy (hba):

\[Average Speed = s/t=720/(\frac{ 17 }{ 3 })\]

OpenStudy (hba):

\[s/t=\frac{720\times3}{ 17 }\]

OpenStudy (hba):

@skittles223 Understood ? now tell me the answer ?

OpenStudy (anonymous):

127.05

OpenStudy (hba):

You Are Wrong :(

OpenStudy (anonymous):

what is it then i cant figure it out im sorry i am so complicated to understand stuff

OpenStudy (hba):

127.05 miles/hour ( approx) - Answer

OpenStudy (hba):

I told you not to forget the units :(

OpenStudy (hba):

You never listen -

OpenStudy (anonymous):

what units

OpenStudy (hba):

Miles/Hour Is The Unit Here -

OpenStudy (hba):

Miles Is the Unit Of S-Distance and Hour is the unit of t-Time

OpenStudy (anonymous):

ok sorry my bad ill work on it

OpenStudy (hba):

Please try to understand it is important -

OpenStudy (anonymous):

Evaluate x3y * x2y2 Answer: a. x to the sixth, y squared b. 5x3y c. 6x2y d. x to the fifth, y cubed e. xy

OpenStudy (hba):

\[ x^3y \times x^2y^2\]

OpenStudy (anonymous):

i cant read that hba

OpenStudy (hba):

Please use magnifier :)

OpenStudy (anonymous):

how do i do that

OpenStudy (hba):

Do you have windows 7 ?

OpenStudy (anonymous):

no im on laptop can u just explain to me wat to do please

OpenStudy (hba):

Windows Is An Operating system in your laptop :(

OpenStudy (anonymous):

why cant u just please tell me how to do it

OpenStudy (hba):

Ok Jus Add The Powers

OpenStudy (hba):

When Multiplying Just Add the powers :)

OpenStudy (anonymous):

like will you type it out for me step by step just dont give me the answer

OpenStudy (hba):

So What are you going to do then ?

OpenStudy (anonymous):

figure out the answer or whats the power to x and y

OpenStudy (hba):

When Multiplying \[x^a \times x^b=x ^{a+b}\]

OpenStudy (anonymous):

does that say Ia times Ib=Ia+b

OpenStudy (hba):

No x^a * x^b = x^{a+b}

OpenStudy (anonymous):

and isnt a 2

OpenStudy (hba):

What Are you trying to say ?

OpenStudy (anonymous):

@skittles223 Can you re-post the current question?